OpenStudy (anonymous):

Which of the following is the fourth term in the expansion of (3 – x)5?

OpenStudy (anonymous):

Its (3-x)^5

OpenStudy (anonymous):

I will love you forever and ever if you help me!

OpenStudy (anonymous):

\[(3-x)^5=-x^5+15 x^4-90 x^3+270 x^2-405 x+243 \]

OpenStudy (apoorvk):

For the expansion of (a-b)^n \[\large T_r = ^nC_r .a^{n-(r-1)}b^{r-1}\]

OpenStudy (anonymous):

Thank you so much, you are my hero!!!!

OpenStudy (apoorvk):

Loolll^